A machine raises a load of 900N by means of an effort of 150N. The distance move by the load is 300mm and the effort moves through a distance of 2300mm in the same time.


a) Calculate the efficiency of the machine at this load. b) Calculate the frictional effort for this load.


1
Expert's answer
2021-08-13T10:20:53-0400

a) The efficiency is true ratio between the output and input work:

"\eta=\frac{W_o}{W_i}=\frac{F_ld_l}{F_ed_e}=\\\space\\=\frac{900\cdot 0.3}{ 150\cdot2.3}=0.78."

b) The frictional effort is


"F_f=F_e-F_l/(d_e/d_l)=32\text N."